# Phase diagram and coarsening dynamics in dry apolar active nematics with reciprocal local alignment

### v2

@@ 4 Mean field and hydrodynamic analysis
is determined by the local scalar order $S({\bf r},t)$ and orientation $\theta({\bf r},t)$. Using active current $J^{(a)}_{i}=-Pe\,\partial_{j}\Pi_{ji}$ [4], the particle density field evolves as,
$\partial_{t}\rho=Pe\,\partial_{i}\partial_{j}\Pi_{ji}+D\nabla^{2}\rho+\partial_{i}c_{i},$ (6)
@@
where $c_{i}$ denotes components of random current taken to be statistically isotropic and delta correlated, and $D$ denotes the diffusion constant. The evolution of nematic order follows,
$\partial_{t}\Pi_{ij}$ $=$ $[\alpha_{1}-\alpha_{2}\,{\rm Tr}(\Pi^{2})]\Pi_{ij}+D\nabla^{2}\Pi_{ij}$ (7)
@@
$+\dfrac{Pe}{4}\left[\partial_{i}\partial_{j}-\frac{1}{2}\delta_{ij}\nabla^{2}\right]\rho+q_{ij}$
where $q_{ij}$ is an uncorrelated Gaussian tensorial noise, which is traceless and symmetric, denoting thermal or active fluctuations [4]. Here, the coefficients $\alpha_{1}=[1-{D_{\theta}}/{D_{\theta}^{(c)}}]$ and $\alpha_{2}=({1}/{4})[{D_{\theta}^{(c)}}/{D_{\theta}}]^{2}$ with the equilibrium mean-field critical point $D_{\theta}^{(c)}=A\rho$ [29]; see Appendix-C.
@@
Figure 2: Typical configurations corresponding to the evolution after a quench from $D_{\theta}=2.8$ to $D_{\theta}=0.28$ at $Pe=428.57$ are shown. The color palette in (a) encodes orientations of particle heading directions. The black arrows show the direction of increasing time. At initial times, $t_{1}=0.007\,D_{r}^{-1}$, the system is in a homogeneous disordered phase (a). By $t_{2}=2.8\,D_{r}^{-1}$, thin nematic bands of higher density are formed. They are oriented in multiple directions (b). At a later time, $t_{3}=21\,D_{r}^{-1}$, these bands coarsen into thicker and longer branched patterns (c). At an even later time $t_{4}=350\,D_{r}^{-1}$, they coarsen into broader and system-spanning bands (d). By the time $t_{5}=12600\,D_{r}^{-1}$, steady-state system-spanning nematic bands appear (e). These bands break, form, and orient along spontaneously chosen directions that vary with time. For example, see another steady-state configuration at $t_{6}=19600\,D_{r}^{-1}$ (f). Here we use $N=64000$, $Pe=428.57$, $\rho r_{c}^{2}=0.4$, and $L=400r_{c}$.
@@
In the homogeneous limit of constant $\rho$, Eq.(7) gives $\frac{dQ_{ij}}{dt}=[\alpha_{1}-\alpha_{2}{\rm Tr}(Q^{2})]Q_{ij}$ leading to a continuous transition at $\alpha_{1}=0$. Describing the first-order transition requires incorporating the density fluctuations [45, 46], for which we use a renormalized mean field theory [47, 32]. We consider a small activity-induced density fluctuation $\delta\rho$ such that $\rho=\rho_{0}+\delta\rho$ to express $\alpha_{1,2}(\rho)=\alpha_{1,2}^{0}+{\alpha^{\prime}}_{1,2}^{0}\delta\rho$ where $\alpha_{1,2}^{0}=\alpha_{1,2}(\rho_{0})$ and ${\alpha^{\prime}}_{1,2}^{0}=(\partial\alpha_{1,2}/\partial\rho)_{\rho_{0}}$. The requirement of a zero current steady state in density evolution suggests $\delta\rho\approx PeS/D$ [29]. Using this in the mean-field limit of Eq.(7), we obtain the evolution for scalar order
@@
where, $u_{2}=\alpha_{1}^{0}$, $u_{3}={\alpha^{\prime}}_{1}^{0}Pe/D$ and $u_{4}=\alpha_{2}^{0}$. This equation can be expressed as a free energy minimizing kinematics $\partial_{t}S=-\partial{\cal F}/\partial S$ with
@@
Figure 3: The nematic and density autocorrelation $C_{\bf Q}(r)$ ($a$) and $C_{\rho}(r)$ ($b$) during quench from $D_{\theta}=2.8$ to $D_{\theta}=0.28$ at $Pe=428.57$. Insets show the correlations plotted against $r/L$. The main plots show data collapse by scaling separation $r$ by dynamic correlation lengths $\xi(t)$ ($a$) and $\zeta(t)$ ($b$). $(c)$ The correlation lengths $\xi(t)$ ($\square$), $\zeta(t)$ ($\circ$) in the scaling regime exhibit similar power-law growths $\sim t^{1/z}$ with exponent $z\approx 2.5$, indicated by the solid green line. The growth exponent $32<z<$ 3 is shown using the dashed blue line of $t^{1/2}$ scaling and the dash-dotted black line of $t^{1/3}$ scaling. Cells of linear dimension $5\,r_{c}$ are used to calculate the coarse-grained correlations from numerical simulations.
@@ 5 Phase ordering kinematics
Having established the phenomenology of the NI transition, we now aim to understand the coarsening kinematics of the nematic order. Towards this end, we present our numerical results of the growth kinetics following a deep quench at a fixed $Pe=428.57$ from a homogeneous isotropic initial state at $D_{\theta}=2.8$ to $D_{\theta}=0.28$ corresponding to a steady-state nematic. For better statistics, we use a larger system with $N=64000$. Fig.2 shows a typical series of snapshots depicting this evolution. As time progresses, domains of higher order parameters and higher density form, with their typical size increasing over time. The deep quench leads to local instability. In the beginning, the homogeneous isotropic state shows instability through the formation of nematically ordered filaments crisscrossing each other; see Fig.2(b). The small filaments merge and coarsen with time, as shown in Fig.2(c) and (d). At a late time, one finds dynamic system-spanning bands of nematically ordered regions; see Fig.2(e) and (f). In the steady state, the local undulations in nematic bands make them unstable toward breaking, and one finds repeated formation, breaking, and reorientations of such bands with time [12, 11].
@@
where $\delta\rho(r)=\rho(r)-\rho_{0}$ and $r$ denotes the separation between local coarse-grained volumes. For further details, see Appendix-D.
@@
In the scaling regime, the correlation lengths grow following power laws $\xi(t)\sim{t}^{1/z_{S}}$ and $\zeta(t)\sim t^{1/z_{\rho}}$, with dynamic exponents $z_{S},z_{\rho}\approx 2.5$; see Fig.3($c$). It is interesting to note that earlier equilibrium simulations of a soft-spin version of the 2D Lebowhl-Lasher model led to an approximate phase ordering exponent $z_{S}\approx 2.5$ [48] differing from $\xi\propto(t/\ln t)^{1/2}$ growth law expected for equilibrium nematic [49]. In our simulations, the value of $z_{\rho}$ is expected to be related to $z_{S}$ due to the density-order coupling. Indeed, the bounds of non-conserved and conserved dynamics [50] to these exponents, $32<z_{S},z_{\rho}<$ 3, are consistent with earlier non-stochastic hydrodynamic estimates [43]. A precise determination of the growth laws and growth exponents requires performing a more careful and extensive finite-size scaling procedure.
@@ 6 Outlook
Moreover, we investigated phase ordering kinetics by quenching from the isotropic to the nematic phase. This revealed coarsening patterns with correlation lengths in nematic order and density fluctuations growing as $t^{1/z}$ with exponents, $z_{S},z_{\rho}\approx 2.5$ remaining within the bounds of non-conserved and conserved dynamics, $32<z_{S},\,z_{\rho}<$ 3, a range consistent with earlier hydrodynamic predictions in Ref. [43]. The quench in scaled orientational coupling showed a lag in particle clustering relative to the growth of nematic correlation.
@@ Acknowledgments
Appendix A Giant number fluctuations
@@
Appendix B Density-nematic order cross correlation
Appendix C Equilibrium mean-field theory for nematic-isotropic transition
@@
Appendix D Coarse-grained correlations

### v3

@@ 4 Mean field and hydrodynamic analysis
is determined by the local scalar order $S({\bf r},t)$ and orientation $\theta({\bf r},t)$.
In the following, we present a mean-field description of the coupled evolution of the density $\rho({\bf r},t)$ and order parameter field $Q({\bf r},t)$ for active apolar nematics [4, 2, 16, 18].
Local curvature in the nematic field can induce a temporary polarity, which, when activity is present, drives a current in the direction determined by this polarity [4]. This results in an active current given by $J^{(a)}_{i}=-Pe\,\partial_{j}\Pi_{ji}$. Using the active current $J^{(a)}_{i}$ and the diffusive current $J^{(d)}_{i}=-D\partial_{i}\rho$, the evolution of particle density field, governed by the continuity equation $\partial_{t}\rho=-\partial_{i}(J^{(a)}_{i}+J^{(d)}_{i})$ leads to,
$\partial_{t}\rho=Pe\,\partial_{i}\partial_{j}\Pi_{ji}+D\nabla^{2}\rho,$ (6)
@@
where $D$ denotes the diffusion constant. The evolution of nematic order follows [4, 2, 18],
@@
$\partial_{t}\Pi_{ij}=[\alpha_{1}-\alpha_{2}\,{\rm Tr}(\Pi^{2})]\Pi_{ij}+D\nabla^{2}\Pi_{ij}+\dfrac{Pe}{4}\left[\partial_{i}\partial_{j}-\frac{1}{2}\delta_{ij}\nabla^{2}\right]\rho\,.$ (7)
In the above equation, the term $[\alpha_{1}-\alpha_{2}\,{\rm Tr}(\Pi^{2})]\Pi_{ij}$ can be derived from a Landau-de Gennes free energy functional, which effectively describes an equilibrium continuous transition between the nematic and isotropic phases in 2D [60, 61]. For a traceless symmetric tensor $\Pi$, the conditions $\text{Tr}(\Pi^{3})=\text{Tr}(\Pi)=0$ in 2D imply that cubic order terms in $\Pi$ are not present in the free energy functional. Consequently, this leads to the absence of quadratic order terms in $\Pi$ in the equation above. As we show in Appendix-D, the coefficients in the Landau-de Gennes theory are given by $\alpha_{1}=[1-{D_{\theta}}/{D_{\theta}^{(c)}}]$ and $\alpha_{2}=({1}/{4})[{D_{\theta}^{(c)}}/{D_{\theta}}]^{2}$ with the mean-field critical point $D_{\theta}^{(c)}=A\rho$. The $D\nabla^{2}\Pi_{ij}$ term in the above equation is the Ginzburg-like term in the traditional Landau-Ginzburg description [61]. The last term in Eq.(7), which couples the evolution of nematic order to density field gradients, arises solely from the active processes [4, 18].
The interplay between the conserved density field and the non-conserved order-parameter field resembles Model C from the Hohenberg-Halperin classification [62]. However, when activity is absent (Pe = 0), the evolution of density and nematic order within our model gets decoupled. Consequently, our phase diagram and the coarsening dynamics in density have no equilibrium counterpart.
@@
Figure 2: Typical configurations corresponding to the evolution after a quench from $D_{\theta}=2.8$ to $D_{\theta}=0.28$ at $Pe=428.57$ are shown. The color palette in (a) encodes orientations of particle heading directions. The black arrows show the direction of increasing time. At initial times, $t_{1}=0.007\,D_{r}^{-1}$, the system is in a homogeneous disordered phase (a). By $t_{2}=2.8\,D_{r}^{-1}$, thin nematic bands of higher density are formed. They are oriented in multiple directions (b). At a later time, $t_{3}=21\,D_{r}^{-1}$, these bands coarsen into thicker and longer branched patterns (c). At an even later time $t_{4}=350\,D_{r}^{-1}$, they coarsen into broader and system-spanning bands (d). By the time $t_{5}=12600\,D_{r}^{-1}$, steady-state system-spanning nematic bands appear (e). These bands break, form, and orient along spontaneously chosen directions that vary with time. For example, see another steady-state configuration at $t_{6}=19600\,D_{r}^{-1}$ (f). Here we use $N=64000$ particles, $Pe=428.57$, $\rho_{0}r_{c}^{2}=0.4$, and $L=400r_{c}$.
@@
Within the mean-field approximation of uniform density and order, Eq.(7) gives $\frac{dQ_{ij}}{dt}=[\alpha_{1}-\alpha_{2}{\rm Tr}(Q^{2})]Q_{ij}$ leading to a continuous transition at $\alpha_{1}=0$, i.e., at $D_{\theta}=D_{\theta}^{(c)}$ Describing the first-order transition in the nematic order requires incorporating the activity induced density fluctuations [63, 64], for which we use a renormalized mean field theory [65, 47]. We consider a small activity-induced density fluctuation $\delta\rho$ such that $\rho=\rho_{0}+\delta\rho$ to express $\alpha_{1,2}(\rho)=\alpha_{1,2}^{0}+{\alpha^{\prime}}_{1,2}^{0}\delta\rho$ where $\alpha_{1,2}^{0}=\alpha_{1,2}(\rho_{0})$ and ${\alpha^{\prime}}_{1,2}^{0}=(\partial\alpha_{1,2}/\partial\rho)_{\rho_{0}}$. The requirement of a zero current steady state in density evolution suggests $\delta\rho\approx Pe\,S/D$ [45]. Using this in the mean-field limit of Eq.(7), we obtain the evolution for scalar order
@@
where, for notational simplicity we used $u_{2}=\alpha_{1}^{0}$, $u_{3}={\alpha^{\prime}}_{1}^{0}Pe/D$ and $u_{4}=\alpha_{2}^{0}$. This equation can be expressed as a free energy minimizing kinematics $\partial_{t}S=-\partial{\cal F}/\partial S$ with
@@
Figure 3: The nematic and density autocorrelation $C_{\bf Q}(r)$ ($a$) and $C_{\rho}(r)$ ($b$) during quench from $D_{\theta}=2.8$ to $D_{\theta}=0.28$ at $Pe=428.57$. Insets show the correlations plotted against $r/L$. The main plots show data collapse by scaling separation $r$ by dynamic correlation lengths $\xi(t)$ ($a$) and $\zeta(t)$ ($b$). $(c)$ The correlation lengths $\xi(t)$ ($\square$) for nematic order, $\zeta(t)$ ($\circ$) for density exhibit similar power-law growths $\sim t^{1/z}$ in the scaling regime with exponents $z_{S}\approx 2.5\pm 0.03$ and $z_{\rho}=2.34\pm 0.02$ (fits are shown using lines through data), values not too different from the numerical estimate of $z_{S}\approx 2.5$ for equilibrium Lebowhl-Lasher model as indicated by the solid maroon line. The model A growth law $t^{1/2}$ (dashed blue line) and the model B growth law $t^{1/3}$ (dash-dotted black line) are shown for comparison. Cells of linear dimension $5\,r_{c}$ are used to calculate the coarse-grained correlations from numerical simulations.
@@
Incorporating spatial variations through the hydrodynamic equations 6 and 7 for apolar nematics helps explain the band formation observed in the coexistence region [18, 32]. Subcritical bifurcations of band solutions in direction-reversing self-propelled rods were demonstrated in Ref. 32. The term $\partial_{i}\partial_{j}\Pi_{ij}$ is crucial for the development of density inhomogeneities [16]. Similar band formations have been noted in polar nematics and analyzed with hydrodynamic equations [31].
@@ 5 Phase ordering kinematics
Having established the phenomenology of the NI transition, we now aim to understand the coarsening kinematics of the nematic order. Towards this end, we present our numerical results of the growth kinetics following a deep quench at a fixed $Pe=428.57$ from a homogeneous isotropic initial state at $D_{\theta}=2.8$ to $D_{\theta}=0.28$ corresponding to a steady-state nematic. For better statistics, we use a larger system with $N=64000$. Fig.2 shows a typical series of snapshots depicting this evolution. As time progresses, domains of higher order parameters and higher density form, with their typical size increasing over time. The deep quench leads to local instability. In the beginning, the homogeneous isotropic state shows instability through the formation of nematically ordered filaments crisscrossing each other; see Fig.2(b). The small filaments merge and coarsen with time, as shown in Fig.2(c) and (d). At a late time, one finds dynamic system-spanning bands of nematically ordered regions; see Fig.2(e) and (f). In the steady state, the local undulations in nematic bands make them unstable toward breaking, and one finds repeated formation, breaking, and reorientations of such bands with time [20, 16, 17].
@@
where $\delta\rho(r)=\rho(r)-\rho_{0}$ and $r$ denotes the separation between local coarse-grained volumes. For further details, see Appendix-E.
@@
In the scaling regime of the deep quench in our active system, the correlation lengths grow following power laws $\xi(t)\sim{t}^{1/z_{S}}$ and $\zeta(t)\sim t^{1/z_{\rho}}$. By fitting the correlation lengths within the scaling regime from $tD_{r}=1$ to $100$, we obtain dynamic exponents $z_{S}=2.5\pm 0.03$ and $z_{\rho}=2.34\pm 0.02$; see Fig.3($c$). Over this scaling regime, correlation functions of nematic order and density show reasonable data collapse as observed in Fig.3($a$) and ($b$). At shorter times, significant correlated domains have not yet formed, while at longer times, finite-size effects become noticeable.
As we noted before, density and order parameters decouple in the passive limit of $Pe=0$. As a result, a quench in $D_{\theta}$ influences nematic order alone. However, even estimating the growth exponent $z_{S}$ in 2D passive nematics is not that simple. For example, a soft-spin variant of the 2D Lebowhl-Lasher model estimated $z_{S}\approx 2.5$ [66], but could not validate the theoretical growth law $\xi\propto(t/\ln t)^{1/2}$ [67]. The active phase ordering exponent $z_{S}$ is consistent with that of the 2D equilibrium Lebwohl-Lasher model [66], and the growth kinetics of density domains, characterized by $z_{\rho}$, surpass the rate predicted for model B dynamics of conserved quantities [61]. Additionally, our findings differ from those of equilibrium model C dynamics, where coupling a conserved field with a non-conserved order parameter of $n>1$ components renders the influence of a non-critical conserved variable unimportant, resulting in a dynamical exponent equivalent to that of model A [62, 61].
A precise determination of the growth laws and growth exponents for stochastic models of active nematic requires performing a more careful and extensive finite-size scaling procedure.
@@ 6 Outlook
Moreover, we investigated phase ordering kinetics by quenching from the isotropic to the nematic phase. This revealed coarsening patterns with correlation lengths in nematic order and density fluctuations growing as $t^{1/z}$ with exponents. The phase ordering exponent $z_{S}=2.5\pm 0.03$ aligns with 2D equilibrium Lebwohl-Lasher model, and the exponent $z_{\rho}=2.34\pm 0.02$ controlling growth kinetics of density domains predicts a faster kinetics than the the rate expected for conserved quantities. Both the exponents $z_{S}$ and $z_{\rho}$ remain within the bounds of model A exponent 2 and model B exponent 3, consistent with earlier hydrodynamic predictions [57]. The quench in scaled orientational coupling showed a lag in particle clustering relative to the growth of nematic correlation. Our predictions are testable in experiments.
@@
Conflicts of interest
There are no conflicts to declare.
Data availability
The data supporting this article have been included as part of the ESI.
@@ Acknowledgments
Appendix A Nematic order parameter in 2D
In two dimensions (2D), the local nematic order can be expressed as,
$\mathbf{Q}({\bf r})=\frac{1}{2}\begin{pmatrix}\langle\cos 2\theta_{i}\rangle&\langle\sin 2\theta_{i}\rangle\\ \langle\sin 2\theta_{i}\rangle&-\langle\cos 2\theta_{i}\rangle\end{pmatrix}$ (14)
[…]
$S({\bf r})=[\langle\cos(2\theta_{i})\rangle^{2}+\langle\sin(2\theta_{i})\rangle^{2}]^{1/2}\,.$ (15)
In the eigenbasis, the order parameter takes the form
$\mathbf{Q}({\bf r})=\frac{1}{2}\begin{pmatrix}S({\bf r})&0\\ 0&-S({\bf r})\end{pmatrix}\,.$ (16)
Note that by definition, the nematic order parameter is bounded by $0\leq S\leq 1$.
Appendix B Giant number fluctuations
@@
Appendix C Density-nematic order cross correlation
@@
Appendix D Equilibrium mean-field theory for nematic-isotropic transition
@@
$(a)\,t=t_{i}$
$(b)\,t_{ii}$
$(c)\,t_{iii}$
Figure 6:  Configurations of nematic bands at $(Pe,D_{\theta})=(714.28,1.166)$: $(a)$ just before the formation of nematic band at time $t_{i}=21315\,D_{r}^{-1}$, $(b)$ when the band is fully formed at $t_{ii}=21490\,D_{r}^{-1}$, and $(c)$ as the band disintegrates at $t_{iii}=21980\,D_{r}^{-1}$. The color palette in (a) encodes orientations of particle heading directions.
@@
Appendix E Coarse-grained correlations
@@
Appendix F Nematic Bands in the coexistence region
In Fig. 6, we present three snapshots corresponding to the parameter values $(Pe,D_{\theta})=(714.28,1.166)$, the parameters in Fig. 1(f) of the main text. The snapshots capture distinct moments during the simulation: $(a)$ just before nematic bands form at some time instance $t_{i}=21315\,D_{r}^{-1}$, $(b)$ when the band is fully formed at $t_{ii}=t_{i}+175\,D_{r}^{-1}$, and $(c)$ as the band disintegrates at $t_{iii}=t_{i}+665\,D_{r}^{-1}$. This cycle of nematic band formation and fragmentation occurs repeatedly, with the band orientations changing spontaneously between such cycles.